Need help, find the domain of the composite function f*g; f(x)=^/x;g(x)=4x+4
Thank you

Answer by stanbon(57387)

find the domain of the composite function f*g; f(x)=^/x;g(x)=4x+4
----------------
f[g(x)] = f[4x+4] = sqrt(4x+4)= 2sqrt(x+1)
Domain: all Real Numbers >=-1
Range: Real Numbers >= 0
=================
Cheers,
Stan H.